pls help

A slope passing through (9,15) and (2,y) is parallel to a line with slope 2. What is the value of y?

Answer by checkley71(8403)

SLOPE=(Y-15)/(2-9)=2 CROSS MULTIPLY
2*-7=Y-15
-14=Y-15
Y=-14+15
Y=1 ANSWER. (9,15)(2,1)
PROOF
(1-15)/2-9)=2
-14/-7=2
2=2